# Learning Contrastive Feature Representations for Facial Action Unit Detection

### v7

@@ 1 Introduction
Facial Action Coding System (FACS) [1] is a widely employed approach for facial expression coding, which introduced a set of 44 facial action units (AUs) that establish a link between facial muscle movements and facial expressions [2]. AU detection is a multi-label binary classification problem. Each AU represents a kind of label, and activated and inactivated AUs are respectively labeled 1 and 0. Nowadays, AU detection [3] has emerged as a CV technique to automatically identify AUs in videos or images. This advancement holds promise for diverse applications, including human-computer interaction, emotion analysis, car driving monitoring et al..
In recent years, supervised learning has dominated AU detection research [4, 5], typically employing binary cross-entropy (BCE) loss for optimization. However, the emphasis of BCE on precise label alignment and pixel-level encoding often leads to overfitting and excessive model complexity, limiting generalization to unseen data. Shifting from pixel-level features to inter-AU differential encoding allows the model to capture AU-specific spatial dependencies by emphasizing relative variations among AUs rather than absolute intensities. This abstraction is theoretically consistent with the Facial Action Coding System (FACS), where facial expressions emerge from the coordinated activation of multiple AUs.
@@
Representation learning based on discriminative contrastive learning has shown substantial potential [6] for classification tasks. As illustrated in Fig. 1, it maximizes similarity among semantically positive samples while minimizing it among negative ones, effectively optimizing mutual information [7]. Applying this framework to AU detection, however, faces two key challenges: severe class imbalance, where inactivated AUs dominate and bias predictions toward the negative class [8], and annotation noise, arising from subtle AU activations that complicate labeling and undermine model fitting [9].
To address these challenges, we propose a discriminative contrastive learning framework inspired by self-supervised learning for learning highly semantic image representations. This framework encodes differential information between AUs from each sample pair, providing an alternative to conventional pixel-level feature learning. The resulting AUNCE loss, conceptually similar to InfoNCE, is complemented by a negative sample re-weighting strategy [10] to mitigate class imbalance by assigning varying importance to individual samples. Additionally, a positive sample sampling strategy incorporates self-supervised signals via defined sampling ratios, ensuring that robust positive samples are derived through data augmentation.
Our research question is whether inter-AU differential encoding, combined with discriminative contrastive learning, can improve AU-specific representation learning and robustness under class imbalance and label noise. We hypothesize that such encoding enables more generalized and semantically consistent AU detection compared to conventional pixel-level feature learning approaches. Unlike prior purely self-supervised frameworks such as SimCLR [11], MoCo [12], KSRL [13], CLEF [14], et al., our approach integrates both supervised and self-supervised signals to balance semantic discrimination and representation robustness. Specifically, supervised AU labels (derived from manual annotations) enable the model to capture semantic category information instead of meaningless instance-level features—addressing a key limitation of purely self-supervised methods, which tend to overemphasize instance-specific differences. Meanwhile, self-supervised consistency signals (mined from data augmentations) enhance the model’s robustness against class imbalance and annotation noise. This joint design provides a principled means of more effectively leveraging limited yet noisy AU labels, thereby distinguishing our approach from existing contrastive-learning methods that lack such task-specific semantic guidance.
The primary contributions of our work are summarized as follows:
@@ 2.1 Discriminative contrastive learning
Contrastive learning [7] originated in self-supervised learning, where tasks are assigned to agents by applying various augmentations to the same image and evaluating feature similarity. The goal is to maximize similarity between features of the same image while minimizing it between different images. Over time, contrastive learning has expanded to supervised and semi-supervised settings, leveraging labeled data to refine feature representations.
In recent years, discriminative contrastive learning, a subset of deep self-supervised learning methods [11, 12], has witnessed substantial progress. This approach follows a learn-to-compare paradigm, distinguishing between similar (positive) and dissimilar (negative) data points in a feature space. For instance, Oord et al.[7] introduced the InfoNCE loss to optimize feature extraction by maximizing consistency in sequence data, establishing a foundation for contrastive methods. Chen et al.[11] proposed SimCLR, which uses a base encoder and projection head to maximize agreement through InfoNCE. Similarly, MoCo [12] leverages a momentum network to maintain a queue of negative samples for efficient learning. While variations in representation encoders and similarity measures exist, the core principle remains: training a representation encoder $f$ by contrasting similar pairs ($x,x^{+}$) against dissimilar pairs ($x,x^{-}$) to optimize the contrastive loss.
In addition to enabling the automatic extraction of fine-grained feature representations via the comparison of positive and negative pairs, deep contrastive learning has evolved to incorporate diverse contrastive objectives tailored for different supervision levels. In this context, Khosla et al. [15] extended the contrastive learning concept to supervised settings with a supervised contrastive learning loss, aiming to enhance feature expression using labeled data. Additionally, Robinson et al. [16] introduced a label-free strategy to sample hard negatives in contrastive learning, enabling tighter class clustering and improved downstream performance with minimal implementation overhead.
@@ 2.2 Contrastive learning for AU detection
Contrastive learning was firstly designed as a pretext task for unsupervised or self-supervised learning and was introduced to the domain of AU detection with limited annotations [17, 14]. For example, Niu et al. [17] introduced a semi-supervised co-training approach that utilized contrastive loss to enforce conditional independence between facial representations generated by two Deep Neural Networks (DNNs). Following this, Zhang et al. [14] introduced a weakly-supervised contrastive learning approach that effectively utilizes coarse-grained activity information.
Subsequent research endeavors extended contrastive learning to supervised settings, aiming at enhancing feature representation for AU detection through the utilization of labeled data [13, 18]. Chang et al. [13] present a knowledge-driven representation learning framework aimed for AU detection, incorporating a contrastive learning module that leverages dissimilarities among facial regions to effectively train the feature encoder. In a similar vein, Suresh et al. [18] proposed a feature-based positive matching contrastive loss, which facilitates learning distances between positives based on the similarity of corresponding AU embeddings.
However, previous methods largely relied on weighted cross-entropy and triplet losses to encode pixel-level facial information, introducing redundant signals that can dilute AU-specific features. To mitigate this, Sun et al. [19] proposed EmoCo, an unsupervised contrastive framework using emotion-based image pairs; however, label inconsistencies and dependence on emotion annotations limit its applicability. Li et al. [9] employed a MoCo-based self-supervised approach for dynamic AU detection, yet its reliance on temporal cues restricts generalization to static images. Further, Li et al. [20] combined image- and feature-level contrastive learning to disentangle AU-relevant and domain-specific features, but performance may degrade under large domain shifts or limited AU diversity.
@@ 3.1 Objectives
In this study, we confront three main challenges: 1. How to design a novel framework that guides the encoder to focus on the differential information between AUs from each sample pair [19], going beyond traditional methods reliant on pixel-level facial features, 2. How to tackle the class imbalance issue across different AU types [8] to improve the overall accuracy of our framework. 3. How to handle noisy and false labels [9] to increase the model’s robustness against misannotations.
To address these challenges, Section 3 provides detailed solutions. In Section 3.2, we propose a discriminative contrastive learning framework that directs the encoder to capture differential features between AUs from each sample pair, thus moving away from conventional pixel-level feature learning. To combat the class imbalance issue, Section 3.3.1 introduces a negative sample re-weighting strategy, drawing inspiration from, which prioritizes underrepresented AUs and facilitates the detection of challenging samples. Besides, Section 3.3.2 presents a positive sample sampling strategy that incorporates augmented supervision signals, which helps the model effectively learn from noisy and false data. Finally, Section 3.4 will introduce the encoder structure we used in our framework.
@@ 3.2 Discriminative contrastive learning framework
For the task of AU detection, which falls under the category of multi-label binary classification, the prevalent approach for training models has been the utilization of cross-entropy loss. This loss measures the disparity between predicted labels and ground truth labels. However, there exists variability in the occurrence rates of different AUs, with some AUs exhibiting low occurrence rates while others display high occurrence rates. Consequently, in contrast to the conventional cross-entropy loss, we employ a weighted cross-entropy loss [4] denoted as ${L}_{wce}$
@@
Though cross-entropy loss is widely used and effective, it has limitations in tasks with class imbalance, complex data distributions, or fine-grained distinctions [21]. It can lead to overfitting and poor generalization, particularly without regularization or techniques like label smoothing. In AU detection, where subtle inter-class differences are critical, cross-entropy loss may restrict the model’s ability to generalize to unseen data. Additionally, the requirement for predicted outputs $\hat{y}$ to satisfy specific conditions for each AU imposes a significant burden on the encoder’s learning capacity.
To address this challenge, we introduce a discriminative contrastive learning framework for AU detection. As shown in Fig.2, the training process comprises two stages: pretraining and linear evaluation. The pretraining stage, preceding linear evaluation, integrates a feature encoder with two key components—a positive sample mining module and a negative sample re-weighting module. The former combines self-supervised and supervised signals to mitigate the impact of noisy labels caused by subtle AU variations, while the latter alleviates class imbalance and enhances hard negative sample mining. Further details are provided in Sections 3.3.1 and 3.3.2.
@@
In the linear evaluation phase, we follow a standard protocol commonly adopted in prior work [7, 11], where a linear classifier is trained atop the frozen encoder using weighted cross-entropy loss (WCE). Notably, this classifier is a single fully connected layer, which serves solely to evaluate representation quality, without influencing feature learning. Thus, the model’s performance primarily reflects the effectiveness of AUNCE in guiding representation learning.
3.3 Learning Strategies
3.3.1 Negative sample re-weighting strategy
To address the second challenge posed by class imbalance, we implement a negative sample re-weighting strategy specifically targeting the anchor data points. This also contributes to hard negative sample mining. Inspired by HCL [16], we apply the following weighting scheme to each negative sample $x_{j}^{-}$:
@@ 3.3.1 Negative sample re-weighting strategy
The key difference between our weighting strategy and that of HCL [16] is the omission of mean normalization in Eq. (3), allowing for more precise class imbalance mitigation in AU detection. We set $\beta\geq 1$ for majority class anchors to amplify minority class gradients, and $\beta\leq 1$ for minority class anchors to suppress majority class influence. By adjusting $\beta$, the model dynamically prioritizes underrepresented classes. Including mean normalization would mix $\beta$ with the normalization term, reducing interpretability and hindering gradient control, making our design critical for effective imbalance-aware learning.
3.3.2 Positive sample sampling strategy
@@
Supervised Signal: Label smoothing is a widely used technique to mitigate label noise by introducing uncertainty through softened target distributions. However, its direct application in contrastive learning is non-trivial due to the absence of explicit labels in loss computation. To address this, we leverage the memory effect of neural networks—where models tend to learn clean samples before overfitting noisy ones. This phenomenon, observed across architectures, implies that positive samples with higher similarity to the anchor are more likely to be reliably learned.
@@
Self-supervised Signal: In addition to label noisy learning, we incorporate self-supervised learning techniques into the training process. By designing these auxiliary tasks carefully, we can encourage the model to learn more robust and invariant representations, which can help mitigate the effects of label noise and improve generalization performance. Specifically, our second method to obtain positive feature involves performing semantically invariant image augmentation on the anchor point $x$:
@@
where $\mathcal{T}(\cdot)$ is a stochastic data augmentation module that applies random transformations to a given data example $x$, generating two correlated views referred to as $x_{s1}^{+}$ and $x_{s2}^{+}$, which we consider as a positive pair. Follow SimCLR [11], we employ a sequential application of three simple augmentations: random cropping and flipping followed by resizing back to original size, random color distortions, and random Gaussian blur.
@@
Unlike data augmentation-based methods such as SimCLR [11], which rely on instance-level contrast using two augmented views of the same sample, our approach constructs positive examples as class prototypes—defined by the mean representation of multiple same-class samples. This design captures more stable and representative class-level semantics, enabling more robust feature comparisons. The resulting positive feature $f(x_{m}^{+})$, as defined in Eq. (10), serves as a virtual class centroid synthesized from M positive samples.
@@ 3.4 Encoder for AU detection
To preserve the fairness of the comparison with conventional pixel-level feature learning methods, our encoder is consistent with the network structure of the first training stage in MEGraph [22], consisting of Swin-transformer backbone and AU Relationship-aware Node Feature Learning module. The input to the encoder consists of an contrastive larning image pair consisting of an anchor and a random sample, and the encoder is optimized by AUNCE only.

### v9

@@ 1 Introduction
Facial Action Coding System (FACS) [rosenberg:et.al:2020] is a widely employed approach for facial expression coding, which introduced a set of 44 facial action units (AUs) that establish a link between facial muscle movements and facial expressions [fathallah2017facial]. AU detection is a multi-label binary classification problem. Each AU represents a kind of label, and activated and inactivated AUs are respectively labeled 1 and 0. Nowadays, AU detection [shao2025facial] has emerged as a computer vision technique to automatically identify AUs in videos or images. This advancement holds promise for diverse applications, including human-computer interaction, emotion analysis, car driving monitoring et al.
In recent years, supervised learning has been the dominant paradigm for developing AU detection methods [shao2026constrained]. A common practice in these methods is to use binary cross-entropy (BCE) loss as the optimization objective. However, BCE loss inherently prioritizes strict label alignment and fine-grained pixel-level feature encoding during training. This tendency can lead to two key limitations: first, it may cause the model to overfit to training data (as it overemphasizes minor label or pixel variations that are irrelevant to AU detection); second, it often increases overall model complexity by forcing the network to learn overly detailed features—even though subtle differences in AU-related patterns are usually sufficient for accurate detection. Together, these issues restrict the model’s ability to generalize to unseen data, hindering its performance in real-world scenarios.
@@
As a result, representation learning approaches based on discriminative contrastive learning have emerged and shown significant potential [assran2023self] for classification tasks. As shown in Fig. 1, this approach aims to maximize the consistency of positive samples that are semantically similar while minimizing the consistency of negative samples that are semantically dissimilar. By optimizing the contrastive loss, the mutual information between variables is either explicitly or implicitly optimized [henaff2020data]. However, the direct integration of contrastive learning into AU detection encounters two principal challenges. The first challenge is the class imbalance issue of each AU type [zhou2020bbn]. Most AU classes manifest a substantial imbalance between majority and minority class samples, with inactivated AUs vastly outnumbering their activated counterparts. This imbalance issue influences the classification decision threshold, potentially resulting in the direct classification of the majority of AUs as the negative class. The second challenge pertains to the noisy and false label problem [li2024contrastive]. Due to the subtle feature differences when AUs activate leading to the inherent difficulty in expertly labeling AUs, existing AU datasets may exhibit a few noisy and false labels, significantly undermining the overall fitting ability of models.
To address these challenges, we propose a discriminative contrastive learning framework inspired by self-supervised learning techniques for learning highly semantic image representations [assran2023self]. This framework encodes differential information between AUs from each sample pair, offering an alternative to the traditional pixel-level feature learning methods commonly employed in AU detection. The resulting loss, termed AUNCE, is conceptually similar to the widely used InfoNCE loss. Moreover, to tackle the class imbalance issue in AU detection, we adopt a negative sample re-weighting strategy inspired by [shang2024facial]. This strategy addresses class imbalance by assigning varying levels of importance to individual samples during training, achieved by adjusting the weights of their loss contributions. Additionally, to mitigate the effects of noisy and false AU labels, inspired by the sample-mixing concept proposed in [zhang2017mixup], we introduce a positive sample sampling strategy. This strategy enhances supervision signals by incorporating self-supervised signals, achieved by defining sampling ratios for different types of positive sample pairs. This ensures that more robust positive samples are derived through data augmentation combined with self-supervised signal. The primary contributions of our work are summarized as follows:
@@ 2.1 Discriminative contrastive learning
Contrastive learning [henaff2020data] originated in self-supervised learning, where tasks are assigned to agents by applying various augmentations to the same image and evaluating feature similarity. The goal is to maximize similarity between features of the same image while minimizing it between different images. Over time, contrastive learning has expanded to supervised and semi-supervised settings, leveraging labeled data to refine feature representations.
In recent years, discriminative contrastive learning, a subset of deep self-supervised learning methods [chen2020simple, he2020momentum], has witnessed substantial progress. This approach follows a learn-to-compare paradigm, distinguishing between similar (positive) and dissimilar (negative) data points in a feature space. For instance, Oord et al.[henaff2020data] introduced the InfoNCE loss to optimize feature extraction by maximizing consistency in sequence data, establishing a foundation for contrastive methods. Chen et al.[chen2020simple] proposed SimCLR, which uses a base encoder and projection head to maximize agreement through InfoNCE. Similarly, MoCo [he2020momentum] leverages a momentum network to maintain a queue of negative samples for efficient learning. While variations in representation encoders and similarity measures exist, the core principle remains: training a representation encoder $f$ by contrasting similar pairs ($x,x^{+}$) against dissimilar pairs ($x,x^{-}$) to optimize the contrastive loss.
In addition to enabling the automatic extraction of fine-grained feature representations via the comparison of positive and negative pairs, deep contrastive learning has evolved to incorporate diverse contrastive objectives tailored for different supervision levels. In this context, Khosla et al. [khosla2020supervised] extended the contrastive learning concept to supervised settings with a supervised contrastive learning loss, aiming to enhance feature expression using labeled data. Additionally, Robinson et al. [robinson2021contrastive] introduced a label-free strategy to sample hard negatives in contrastive learning, enabling tighter class clustering and improved downstream performance with minimal implementation overhead.
@@ 2.2 Contrastive learning for AU detection
Contrastive learning was firstly designed as a pretext task for unsupervised or self-supervised learning and was introduced to the domain of AU detection with limited annotations [niu2019multi, zhang2023weakly]. For example, Niu et al. [niu2019multi] introduced a semi-supervised co-training approach that utilized contrastive loss to enforce conditional independence between facial representations generated by two Deep Neural Networks (DNNs). Following this, Zhang et al. [zhang2023weakly] introduced a weakly-supervised contrastive learning approach that effectively utilizes coarse-grained activity information.
Subsequent research endeavors extended contrastive learning to supervised settings, aiming at enhancing feature representation for AU detection through the utilization of labeled data [chang2022knowledge, suresh2022using]. Chang et al. [chang2022knowledge] present a knowledge-driven representation learning framework aimed for AU detection, incorporating a contrastive learning module that leverages dissimilarities among facial regions to effectively train the feature encoder. In a similar vein, Suresh et al. [suresh2022using] proposed a feature-based positive matching contrastive loss, which facilitates learning distances between positives based on the similarity of corresponding AU embeddings.
However, previous methods primarily employed a combination of weighted cross-entropy loss and triplet loss to constrain the training process, encoding pixel-level information of the entire face associated with AUs. This results in significant redundant information that may dilute the discriminative signals necessary for accurate AU detectionTo address this, Sun et al. [sun2021emotion] proposed EmoCo, an unsupervised contrastive learning framework that leverages emotion-based image pairs. Although EmoCo improves AU detection performance, it suffers from label inconsistencies within emotion categories, leading to false positives and negatives, and its dependence on emotion annotations restricts its applicability to datasets lacking such labels. Similarly, Li et al. [li2024contrastive] introduced a self-supervised contrastive approach based on the MoCo framework [he2020momentum] for dynamic AU detection in videos. However, its reliance on temporal cues limits its generalization to static images. To further address cross-domain generalization, Li et al. [li2025decoupled], which combines image- and feature-level contrastive learning to disentangle AU-relevant and domain-specific representations. Despite its effectiveness, its performance may decline under large domain shifts or limited AU data diversity.
@@ 3.1 Objectives
In this study, we confront three main challenges: 1. How to design a novel framework that guides the encoder to focus on the differential information between AUs from each sample pair [sun2021emotion], going beyond traditional methods reliant on pixel-level facial features, 2. How to tackle the class imbalance issue across different AU types [zhou2020bbn] to improve the overall accuracy of our framework. 3. How to handle noisy and false labels [li2024contrastive] to increase the model’s robustness against misannotations.
To address these challenges, Section 3 provides detailed solutions. In Section 3.2, we propose a discriminative contrastive learning framework that directs the encoder to capture differential features between AUs from each sample pair, thus moving away from conventional pixel-level feature learning. To combat the class imbalance issue, Section 3.3 introduces a negative sample re-weighting strategy, drawing inspiration from, which prioritizes underrepresented AUs and facilitates the detection of challenging samples. Besides, Section 3.4 presents a positive sample sampling strategy that incorporates augmented supervision signals, which helps the model effectively learn from noisy and false data. Finally, Section 3.5 will introduce the encoder structure we used in our framework.
@@ 3.2 Discriminative contrastive learning framework
For the task of AU detection, which falls under the category of multi-label binary classification, the prevalent approach for training models has been the utilization of cross-entropy loss. This loss measures the disparity between predicted labels and ground truth labels. However, there exists variability in the occurrence rates of different AUs, with some AUs exhibiting low occurrence rates while others display high occurrence rates. Consequently, in contrast to the conventional cross-entropy loss, we employ a weighted cross-entropy loss [shao2021jaa] denoted as ${L}_{wce}$
@@
Though cross-entropy loss is widely used and effective, it has limitations in tasks with class imbalance, complex data distributions, or fine-grained distinctions [wei2021fine]. It can lead to overfitting and poor generalization, particularly without regularization or techniques like label smoothing. In AU detection, where subtle inter-class differences are critical, cross-entropy loss may restrict the model’s ability to generalize to unseen data. Additionally, the requirement for predicted outputs $\hat{y}$ to satisfy specific conditions for each AU imposes a significant burden on the encoder’s learning capacity.
To address this challenge, we introduce a discriminative contrastive learning framework for AU detection. As shown in Fig.2, the training process comprises two stages: pretraining and linear evaluation. The pretraining stage, preceding linear evaluation, integrates a feature encoder with two key components—a positive sample mining module and a negative sample re-weighting module. The former combines self-supervised and supervised signals to mitigate the impact of noisy labels caused by subtle AU variations, while the latter alleviates class imbalance and enhances hard negative sample mining. Further details are provided in Sections 3.3 and 3.4.
@@
In the linear evaluation phase, we follow a standard protocol commonly adopted in prior work [henaff2020data, chen2020simple], where a linear classifier is trained atop the frozen encoder using weighted cross-entropy loss (WCE). Notably, this classifier is a single fully connected layer, which serves solely to evaluate representation quality, without influencing feature learning. Thus, the model’s performance primarily reflects the effectiveness of AUNCE in guiding representation learning.
3.3 Negative sample re-weighting strategy
To address the second challenge posed by class imbalance, we implement a negative sample re-weighting strategy specifically targeting the anchor data points. This also contributes to hard negative sample mining. Inspired by HCL [robinson2021contrastive], we apply the following weighting scheme to each negative sample $x_{j}^{-}$:
@@ 3.3 Negative sample re-weighting strategy
The primary distinction between our weighting strategy and that of HCL [robinson2021contrastive] lies in the deliberate omission of mean normalization in Eq. (3), which enables more precise and adaptive mitigation of class imbalance in AU detection. Specifically, we set $\beta\geq 1$ when the anchor sample belongs to the majority class, amplifying the gradient contributions of minority class samples. Conversely, when the anchor is from the minority class, $\beta\leq 1$ is used to suppress the influence of majority class samples. By tuning the weighting factor $\beta$, the model dynamically adjusts the relative importance of positive and negative samples, thereby enhancing its focus on underrepresented classes. In contrast, the inclusion of mean normalization would entangle the effect of $\beta$ with the normalization term, compromising interpretability and hindering precise gradient modulation. This design choice is therefore critical to achieving effective imbalance-aware learning.
3.4 Positive sample sampling strategy
@@
3.4.1 Supervised Signal
Label smoothing is a widely used technique to mitigate label noise by introducing uncertainty through softened target distributions. However, its direct application in contrastive learning is non-trivial due to the absence of explicit labels in loss computation. To address this, we leverage the memory effect of neural networks—where models tend to learn clean samples before overfitting noisy ones. This phenomenon, observed across architectures, implies that positive samples with higher similarity to the anchor are more likely to be reliably learned.
@@ 3.4.1 Supervised Signal
3.4.2 Self-supervised Signal
In addition to label noisy learning, we incorporate self-supervised learning techniques into the training process. By designing these auxiliary tasks carefully, we can encourage the model to learn more robust and invariant representations, which can help mitigate the effects of label noise and improve generalization performance. Specifically, our second method to obtain positive feature involves performing semantically invariant image augmentation on the anchor point $x$:
@@ 3.4.2 Self-supervised Signal
where $\mathcal{T}(\cdot)$ is a stochastic data augmentation module that applies random transformations to a given data example $x$, generating two correlated views referred to as $x_{s1}^{+}$ and $x_{s2}^{+}$, which we consider as a positive pair. Follow SimCLR [chen2020simple], we employ a sequential application of three simple augmentations: random cropping and flipping followed by resizing back to original size, random color distortions, and random Gaussian blur.
@@
Unlike data augmentation-based methods such as SimCLR [chen2020simple], which rely on instance-level contrast using two augmented views of the same sample, our approach constructs positive examples as class prototypes—defined by the mean representation of multiple same-class samples. This design captures more stable and representative class-level semantics, enabling more robust feature comparisons. The resulting positive feature $f(x_{m}^{+})$, as defined in Eq. (10), serves as a virtual class centroid synthesized from M positive samples.
@@ 3.5 Encoder for AU detection
To preserve the fairness of the comparison with conventional pixel-level feature learning methods, our encoder is consistent with the network structure of the first training stage in MEGraph [luo2022learning], consisting of Swin-transformer backbone and AU Relationship-aware Node Feature Learning module. The input to the encoder consists of an contrastive larning image pair consisting of an anchor and a random sample, and the encoder is optimized by AUNCE only.